Aardvark Industries is considering a project that will generate the following free cash flows:

Year

0

1

2

3

Free Cash Flows

($200)

$100

$80

$60

You are also provided with the following market value balance sheet and information regarding Aardvark's cost of capital:

Assets

Liabilities

Cost of Capital

Cash

0

Debt

400

Debt

7%

Other Assets

1000

Equity

600

Equity

12%

τc

35%

A. Aardvark's unlevered cost of equity is closest to:

A) 10.0% B) 10.4% C) 9.5% D) 9.0%

B. The unlevered value of Aardvark's new project is closest to:

A) $205 B) $100 C) $164 D) $202

C. Suppose that to fund this new project, Aardvark borrows $120 with the principal to be paid in three equal installments at the end each year. The present value of Aardvark's interest tax shield is closest to:

A) $5.15 B) $5.00 C) $5.90 D) $5.26
Answer #1

a) I think there is some problem with this, it should be un-levered cost of capital (in spite of equity).

Unlevered Cost = (D/(D+E)) * (1 -T)* (rd) + (E/(E+D)) * (re) = pre tax WACC ( as this project has the similar risk to Aardvark's other investment , Hence its unlevered cost of capital would be simialr to organization)

=(400/1000) * (7) +(600/1000)* (12)

= 10% (Option A is correct)

b) Unlevered Value = Present Value of Future CashFlow

=($100/1.1^1) +( $80/ 1.1^2) + $60/ (1.1^3)

= $ 202.103 (Option D is correct)

c) Present value of interest shield = Present value of Debt * Interest Rate * Taxrate discounted by interest rate.

(120 * 0.07* 0.35)/1.07^1 + 80*0.07*0.35/1.07^2 + 40*0.07*0.35 / 1.07^3

= $5.2595 (Option D is correct)